# Vega: Learning to Drive with Natural Language Instructions

### 26 Mar 2026 v1

@@ Abstract
Vision-language-action models have reshaped autonomous driving to incorporate languages into the decision-making process. However, most existing pipelines only utilize the language modality for scene descriptions or reasoning and lack the flexibility to follow diverse user instructions for personalized driving. To address this, we first construct a large-scale driving dataset (InstructScene) containing around 100,000 scenes annotated with diverse driving instructions with the corresponding trajectories. We then propose a unified vision-language-world-action model, Vega, for instruction-based generation and planning. We employ the autoregressive paradigm to process visual inputs (vision) and language instructions (language) and the diffusion paradigm to generate future predictions (world modeling) and trajectories (action). We perform joint attention to enable interactions between the modalities and use individual projection layers for different modalities for more capabilities. Extensive experiments demonstrate that our method not only achieves superior planning performance but also exhibits strong instruction-following abilities, paving the way for more intelligent and personalized driving systems. Code is available at https://github.com/zuosc19/Vega.
@@ 3.4 Training and Inference
We implement a single-stage training paradigm to cover both action planning and world modeling. For action planning, we train the model to predict the action plan $A_{t}^{(N)}=[A_{t},...,A_{t+N-1}]$ based on past observations $I_{t}^{(-T)}=[I_{t-T},...,I_{t}]$ and current driving instruction $L_{t}$. For world modeling, we train the model to predict future image observation $I_{t+K}$ based on past images $I_{t}^{(-T)}$, current driving instruction $L_{t}$ and action plan $A_{t}^{(N)}$. We use the MSE of the normalized relative action $A$ as action loss:
@@
In the training stage, we optimize a joint objective with loss $\mathcal{L}_{pretrain}=\lambda_{A}\cdot\mathcal{L}_{A}+\lambda_{V}\cdot\mathcal{L}_{V}$. This allows the model to learn world knowledge alongside planning capabilities. In the inference stage, we use Classifier-Free Guidance Diffusion [17] to generate actions, with both image guidance and text guidance enabled. While we primarily focus on the action planning task during inference, the model retains its image generation capabilities from the training stage.
@@ 4.1 Datasets and Benchmarks
  • NAVSIM v1 [5] filters OpenScene to remove near-trivial and erroneous scenes, reducing the train split size to 85k. During evaluation, NAVSIM v1 runs a non-reactive simulation at 10Hz for future 4 seconds, then scores the driving agent with metrics including No at-fault Collision (NC), Drivable Area Compliance (DAC), Time To Collision (TTC), Comfort (Comf.), and Ego Progress (EP). These metrics are aggregated into the Predictive Driver Model Score (PDMS). We use the train split for finetuning and the test split for evaluation.
@@
Table 2: Comparison with state-of-the-art methods on the NAVSIM v1. NC: no at-fault collision. DAC: drivable area compliance. TTC: time-to-collision. C.: comfort. EP: ego progress. PDMS: the predictive driver model score. Abbreviations: 1x Cam (single front-view camera), Nx Cam (surround-view cameras), L (LiDAR). †: Using the best-of-N (N=6) strategy following [85].
@@ 4.2 Implementation Details
Since VLMs struggle to accurately perceive ego-vehicle motion, we generate supplementary rule-based instructions. We classify scenes using speed, acceleration, and turn rate thresholds, converting them into natural language. While these closed-set instructions lack diversity, they provide precise ego-motion cues. We then use them as auxiliary prompts for the VLM, combining both to generate accurate and diverse driving instructions. With this pipeline, we annotated $85109$ scenes from the NAVSIM train split and $12144$ scenes from the NAVSIM test split.
@@ 4.3 Main Results
Figure 4: Ablation of interleaving image-action sequences. We compare the finetuning losses of models trained on non-interleaving sequences (original) and interleaving sequences of different lengths.
@@ 4.4 Experimental Analysis
Future Frame Prediction. During pretraining, the model is trained to predict the first future frame. We ablate future frame prediction with two additional configurations. First, we randomly sample one of the 8 future frames and specify the chosen index in the text prompt. Second, we remove the future frame prediction task altogether. All three configurations are trained on 8 H20 GPUs, with other hyperparameters identical to pretraining in Section 4.2. The results, shown in Table 3, indicate that the task of future frame prediction indeed improves the planning capabilities of the model, but the exact choice of future frame has limited impact on performance.
Interleaving Observation and Action. In our original design, only past images are provided to the model as reference. We argue that interleaving image and action helps the model learn their dynamics, resulting in faster convergence and lower loss during training. Following [33], we pretrain the model with interleaving image-action sequences. We ablate the pretraining image-action sequence length with 2, 4 and 6. During finetuning, we interleave the original 4 past images with 3 past actions. Figure 4 reveals that although models pretrained on interleaving sequences suffer from higher loss in the initial stages of finetuning, which can be attributed to the discrepancy between their pretraining and finetuning designs, they converge significantly faster than models without interleaving sequences, eventually surpassing the latter. In addition, although all interleaving models use the same sequence length during finetuning, those with longer image-action sequences during pretraining show lower losses.
@@
Figure 6: Future image generation conditioned on instructions and actions. In the same scenario, given two sets of instructions, the model plans two action sequences and generates their respective future images. Both action sequences follow their instructions and both images are consistent with their actions.
Visualizations. In addition to Figure 1, we provide two more examples to visualize the effect of instructions on Vega’s ability to adjust the vehicle’s speed according to user instructions in Figure 5. We test two instructions in each scene and plot the predicted trajectories for the next 8 frames on the front-view-camera image as well as the Bird’s-Eye-View (BEV) map. In both scenes, our model successfully increased, decreased, or maintained speed to follow the instructions. We also offer a qualitative evaluation of our pretrained model’s ability to align both its action planning and image generation with user instructions in Figure 6. We select critical scenes where there can be multiple possible courses of action, e.g. approaching the intersection, encountering another vehicle. For each scene, we give the model different sets of instructions, then generate future actions and images sequentially. We observe that Vega is able to generate future actions and images that are consistent with the instructions, indicating that our world modeling framework has successfully helped the model learn the dynamics of the driving environment.

### 30 Mar 2026 v2

@@ Abstract
Vision-language-action models have reshaped autonomous driving by incorporating natural language into the decision-making process. However, most existing pipelines only utilize the language modality for scene descriptions or reasoning and lack the flexibility to follow diverse user instructions for personalized driving. To address this, we first construct a large-scale driving dataset (InstructScene) containing around 100,000 scenes annotated with diverse driving instructions with the corresponding trajectories. We then propose a unified vision-language-world-action model, Vega, for instruction-based generation and planning. We employ the autoregressive paradigm to process visual inputs (vision) and language instructions (language) and the diffusion paradigm to generate future predictions (world modeling) and trajectories (action). We perform joint attention to enable interactions between the modalities and use individual projection layers for different modalities for more capabilities. Extensive experiments demonstrate that our method not only achieves superior planning performance but also exhibits strong instruction-following abilities, paving the way for more intelligent and personalized driving systems. Code is available at https://github.com/zuosc19/Vega.
@@ 3.4 Training and Inference
We implement a single-stage training paradigm to cover both action planning and world modeling. For action planning, we train the model to predict the action plan $A_{t}^{(N)}=[A_{t},...,A_{t+N-1}]$ based on past observations $I_{t}^{(-T)}=[I_{t-T},...,I_{t}]$ and current driving instruction $L_{t}$. For world modeling, we train the model to predict future image observation $I_{t+N}$ based on past images $I_{t}^{(-T)}$, current driving instruction $L_{t}$ and action plan $A_{t}^{(N)}$. We use the MSE of the normalized relative action AA as action loss:
@@
In the training stage, we optimize a joint objective with loss $\mathcal{L}=\lambda_{A}\cdot\mathcal{L}_{A}+\lambda_{V}\cdot\mathcal{L}_{V}$. This allows the model to learn world knowledge alongside planning capabilities. In the inference stage, we use Classifier-Free Guidance Diffusion [17] to generate actions, with both image guidance and text guidance enabled. While we primarily focus on the action planning task during inference, the model retains its image generation capabilities from the training stage.
@@ 4.1 Datasets and Benchmarks
  • NAVSIM v1 [5] filters OpenScene to remove near-trivial and erroneous scenes, reducing the train split size to 85k. During evaluation, NAVSIM v1 runs a non-reactive simulation at 10Hz for 4 seconds, then scores the driving agent with metrics including No at-fault Collision (NC), Drivable Area Compliance (DAC), Time To Collision (TTC), Comfort (Comf.), and Ego Progress (EP). These metrics are aggregated into the Predictive Driver Model Score (PDMS). We use the train split for training and the test split for evaluation.
@@
Table 2: Comparison with state-of-the-art methods on the NAVSIM v1. NC: No at-fault Collision. DAC: Drivable Area Compliance. TTC: Time-To-Collision. C.: Comfort. EP: Ego Progress. PDMS: Predictive Driver Model Score. Abbreviations: 1x Cam (single front-view camera), Nx Cam (surround-view cameras), L (LiDAR). †: Using the best-of-N (N=6) strategy following [85].
@@ 4.2 Implementation Details
Since VLMs struggle to accurately perceive ego-vehicle motion, we generate supplementary rule-based instructions. We classify scenes using speed, acceleration, and turn rate thresholds, converting them into natural language. While these closed-set instructions lack diversity, they provide precise ego-motion cues. We then use them as auxiliary prompts for the VLM, combining both to generate accurate and diverse driving instructions. With this pipeline, we annotated 8510985109 scenes from the NAVSIM train split and 1214412144 scenes from the NAVSIM test split.
@@ 4.3 Main Results
Figure 4: Ablation of interleaving image-action sequences. We compare the training losses of models trained on non-interleaving sequences (original) and interleaving sequences of different lengths.
@@ 4.4 Experimental Analysis
Future Frame Prediction. We ablate future frame prediction with three configurations. First, we always predict the next future frame. Second, we randomly sample one of the 8 future frames and specify the chosen index in the text prompt. Third, we remove the future frame prediction task altogether. All three configurations are trained on 8 H20 GPUs, with other hyperparameters identical to training in Section 4.2. The results, shown in Table 3, indicate that the task of future frame prediction indeed improves the planning capabilities of the model, but the exact choice of future frame has limited impact on performance.
Interleaving Observation and Action. In our original design, only past images are provided to the model as reference. We argue that interleaving image and action helps the model learn their dynamics, resulting in faster convergence and lower loss during training. Following [33], we train the model with interleaving image-action sequences. We ablate the training image-action sequence length with 2, 4 and 6. During training, we interleave the original 4 past images with 3 past actions. Figure 4 reveals that although models trained on interleaving sequences suffer from higher loss in the initial stages of training, which can be attributed to the discrepancy between their training designs, they converge significantly faster than models without interleaving sequences, eventually surpassing the latter. In addition, although all interleaving models use the same sequence length during training, those with longer image-action sequences during training show lower losses.
@@
Figure 6: Future image generation conditioned on instructions and actions. In the same scenario, given three sets of instructions, the model plans three action sequences and generates their respective future images. All action sequences follow their instructions and all images are consistent with their actions.
Visualizations. In addition to Figure 1, we provide two more examples to visualize the effect of instructions on Vega’s ability to adjust the vehicle’s speed according to user instructions in Figure 5. We test two instructions in each scene and plot the predicted trajectories for the next 8 frames on the front-view-camera image as well as the Bird’s-Eye-View (BEV) map. In both scenes, our model successfully increased, decreased, or maintained speed to follow the instructions. We also offer a qualitative evaluation of our model’s ability to align both its action planning and image generation with user instructions in Figure 6. We select critical scenes where there can be multiple possible courses of action, e.g. approaching the intersection, encountering another vehicle. For each scene, we give the model different sets of instructions, then generate future actions and images sequentially. We observe that Vega is able to generate future actions and images that are consistent with the instructions, indicating that our world modeling framework has successfully helped the model learn the dynamics of the driving environment.